I am having trouble with this equation:
5/ (3 + i) is equivalent to what in a + bi form.

Answer by nerdybill(7384)

5/ (3 + i) is equivalent to what in a + bi form.
.
Multiply by (3-i)/(3-i):
5/(3+i) * (3-i)/(3-i)
= 5(3-i)/(3+i)(3-i)
= 5(3-i)/(9-3i+3i+1)
= 5(3-i)/(9+1)
= 5(3-i)/10
= (15-5i)/10
= 15/10 - 5i/10
= (3/2) - (1/2)i (this is the form they want)